1. Placing a 0.5 ml sample of bacteria into a 4.5 ml sterile water, then transferring a 0.5 ml sample from that dilution to aPlease help me solve these problems. I don't understand
Answer #1

1. The dilution of the tube can be determined by dividing the amount of sample by the total volume of the solution = 0.5/5 = 0.1. Again 0.5 ml if this sample was added to the total volume was 5ml. So the dilution would be 0.1 * 0.1 = 0.01. So the dilution is 1:100.

2. The number of bacteria in the sample is expressed as colony forming units per ml (CFU/ml). Here, two.plates are given so the number of bacteria would be the mean of the two = 130+150/2 = 140

The CFU/ml can be calculated using the formula:
cfu/ml = (no. of colonies x dilution factor) / volume of culture plate.

So for at 10-7 dilution, Bacteria/ml = (140)*(10^7) / 0.1 = 1.4 * 1010

The other dilutions show sparce growth so they are not considered.